### v4

# Distribution Matching Distillation Meets Reinforcement Learning

Abstract
Distribution Matching Distillation (DMD) facilitates efficient inference by distilling multi-step diffusion models into few-step variants. Concurrently, Reinforcement Learning (RL) has emerged as a vital tool for aligning generative models with human preferences. While both represent critical post-training stages for large-scale diffusion models, existing studies typically treat them as independent, sequential processes, leaving a systematic framework for their unification largely unexplored. In this work, we demonstrate that jointly optimizing these two objectives yields mutual benefits: RL enables more preference-aware and controllable distillation rather than uniformly compressing the full data distribution, while DMD serves as an effective regularizer to mitigate reward hacking during RL training. Building on these insights, we propose DMDR, a unified framework that incorporates Reward-Tilted Distribution Matching optimization alongside two dynamic distillation training strategies in the initial stage, followed by the joint DMD and RL optimization in the second stage. Extensive experiments demonstrate that DMDR achieves state-of-the-art visual quality and prompt adherence among few-step generation methods, even surpassing the performance of its multi-step teacher model.
1 Introduction
Diffusion models pre-trained on large-scale datasets have achieved remarkable performance in visual generation tasks [dit, sd3, flux, zimage, sdxl]. However, their sampling process typically requires numerous iterative denoising steps [ddim, ddpm, flow-matching], each involving a full forward pass through a large neural network. This requirement makes high-resolution text-to-image synthesis computationally demanding. Furthermore, base models often fail to produce images that align closely with human preferences and aesthetic criteria [pickscore, hpsv2, hpsv3].
To enhance practical utility, researchers typically employ two post-training stages. First, step-distillation [diffinstruct, dmd, lcm, hypersd, dmd2, flashdiff], which compresses the original model into a generator capable of few-step sampling (e.g., 4 steps) for efficient sampling. Among such methods, Distribution Matching Distillation (DMD) [dmd] is widely recognized for its effectiveness in large-scale scenarios [dmd2, dmdx, twinflow, senseflow] and has been adopted in prominent industrial applications [seedream4, zimage, flux-2]. Second, reinforcement learning (RL) [refl, flow-dpo, flowgrpo, ddpo, dancegrpo, srpo, diffusionnft, diffusiondpo] aligns the model with human preferences to improve aesthetic quality. Despite progress in both domains, most existing studies address these objectives in isolation. Early efforts to incorporate RL into the distillation process were largely confined to a two-stage paradigm [diffinstruct++, hypersd, pso, hypernoise] (exploring how to conduct RL on an already distilled model). Consequently, there remains a critical gap in the field: a systematic framework to explore the unification and potential synergy between these two “temporarily separated” stages.
To bridge this gap, we propose DMDR, a unified framework that seamlessly integrates DMD and RL into a synergistic training pipeline. Our core insight is that concurrent optimization yields mutual benefits: (1) Preference-Aware Distillation: RL steers the distillation process toward high-reward regions, breaking the ”performance ceiling” inherent in purely mimicking a teacher; (2) Distributional Regularization: The DMD objective serves as a robust regularizer for RL, mitigating reward hacking by maintaining proximity to the teacher’s manifold. Specifically, DMDR operates in two stages. In the initial stage, we implement Reward-Tilted Distribution Matching (RT-DM), which distills toward a reward-tilted teacher distribution to prioritize human-preferred regions. Concurrently, we introduce Dynamic Distribution Guidance (DynaDG) and Dynamic Renoise Sampling (DynaRS) to stabilize the ”cold-start” phase through annealed distributional overlap maximization. In the second stage, we transition to a joint optimization paradigm. In this phase, the RL objective drives aggressive preference alignment, while the DMD mechanism serves as a robust regularizer to prevent reward hacking.
Our experimental results show that our method leads to state-of-the-art few-step generative models. Moreover, our method is not only compatible with different models (e.g., flow-based, denoising-based) but also with various RL algorithms (e.g., ReFL, DPO, GRPO). This ensures the long-term effectiveness of our method as the multi-step model and RL algorithm evolve.
In summary, our main contributions are as follows:
  • We propose DMDR, which shows that DMD and RL can be trained simultaneously with mutual benefits.
  • We design a Reward-Tilted Distribution Matching and two dynamic distillation training strategies for better integrating RL and distillation in the initial phase.
  • We validate our method on various models and RL algorithms, all of which achieve excellent performance.
2 Method
2.1 Preliminary
Distribution matching distillation. DMD [dmd] compresses a multi-step diffusion model (teacher) into a few-step generator (student) $G$ by minimizing the time-averaged approximate Kullback-Leibler (KL) divergence between the real distribution $p_{\text{real},t}$ and the synthetic distribution $p_{\text{fake},t}$. Note that $G$ is optimized via gradient descent, and this gradient admits a compact expression as the difference of two score functions:
$\nabla_{\!\theta}\,\mathcal{L}_{\text{dmd}}$ $=\mathbb{E}_{t}\!\left[\nabla_{\!\theta}\operatorname{KL}(p_{\text{fake},t}\|p_{\text{real},t})\right]$ (1)
$=-\mathbb{E}_{t}\!\left[\int\!\Bigl(s_{\text{real}}\bigl(F_{t})\bigr)-s_{\text{fake}}\bigl(F_{t}\bigr)\Bigr)\frac{\mathrm{d}G_{\theta}(z)}{\mathrm{d}\theta}\,\mathrm{d}z\right]\!,$
where $z\sim\mathcal{N}(0,\mathbf{I})$, $\theta$ denotes the parameters of $G$, and $F_{t}$ is the forward diffusion operator that injects noise at time $t$ to $G_{\theta}(z)$. The quantities $s_{\text{real}}$ and $s_{\text{fake}}$ are the score functions estimated by diffusion models $\mu_{\text{real}}$ and $\mu_{\text{fake}}$, respectively. During training, the $\mu_{\text{fake}}$ is updated via diffusion loss $\mathcal{L}_{\text{diff}}$ (for denoising-based models, the prediction target is noise, while for flow-based ones, it is velocity) on synthetic samples produced by the few-step generator. Moreover, as introduced in DMD2 [dmd2], $G_{\theta}(z)$ is a noisy synthetic image produced by the current $G$ running several steps, which is called backward simulation. The generator $G$ then denoises these simulated images, and the outputs are supervised with the above loss functions.
Reinforce learning for diffusion models. Although there are various RL algorithms [refl, flowgrpo, diffusiondpo], All their goal is to optimize the diffusion model to generate the samples that can maximize a score given by reward models [hpsv2, pickscore, clipscore]. This score can be expressed as aesthetic, texture, prompt following, and so on. At the same time, a regularization term is often added to balance the trade-off between reward maximization and the deviation from the original model [rlhf]. Thus, the loss can be formulated as follows:
$\mathcal{L}_{\text{rl}}=\mathbb{E}\left[r(\bm{x}_{0})-\operatorname{KL}(p_{new}\|p_{\text{ref}})\right],$ (2)
where $r$ is the reward function, $p_{new}$ and $p_{ref}$ is the distribution of the trainable model and the reference (this can be a pre-train data distribution [refl] or the original model’s [flowgrpo, diffusiondpo, diffusionnft]), $\bm{x}_{0}$ is the clean image that is generated by the training generation model.
2.2 Problem Formulation and Pipeline Overview.
It is noted that DMD effectively reduces the sampling steps of the diffusion model, while RL enables the diffusion model to generate instruction-aligned and human-preferred images. Both of these are crucial for the practical application of the diffusion model. However, at present, the majority of the work [dmdx, dmd2, magicdistillation, flowgrpo, diffusionnft, dancegrpo] only focuses on studying and improving one aspect, without exploring the synergy between the two. And only a small portion of the works [pso, flash-dmd, diffinstruct++, hypersd, hypernoise] attempt to introduce RL into the distillation process, but they still focus on how to conduct RL on a distilled model. This raises a question: Can these two be combined into one stage and benefit each other?
We address this challenge question by proposing DMDR, as shown in Figure 1, the training process of DMDR is divided into two operational stages: Stage 1: RT-DM with Dynamic Distillation. (Section 2.4 and Section 2.5) Stage 2: Joint RL + DMD Optimization(Section 2.3). The following sections will provide the detailed analysis and illustration.
Figure 1: Overview of DMDR. It follows a two-stage training paradigm: Stage 1 (Reward-Tilted Distribution Matching) employs reward-weighted distillation to integrate preference signals into the early distillation phase, stabilized by Dynamic Distribution Guidance (DynaDG) and Dynamic Renoise Sampling (DynaRS). Stage 2 (Joint RL + DMD) performs direct reward maximization regularized by the DMD loss.
2.3 Bring DMD and RL Together with Benefit
We now show formally and empirically that DMD operates as a superior regularization mechanism that subsumes the traditional KL penalty in RL for few-step model, yielding robust joint optimization.
Figure 2: Visualization of the difference of optimization directions.
From the perspective of RL, standard preference optimization requires a penalty (e.g., $\operatorname{KL}(p_{new}\|p_{\text{ref}})$) to prevent reward hacking [refl, flowgrpo]. In few-step models, sequentially applying RL after distillation uses the already-distilled distribution $P_{\text{fake-ref}}$ as the regularizer (Fig. 2). Because distillation inherently loses some modes, $P_{\text{fake-ref}}$ acts as an impoverished anchor, leading to rapid overfitting and mode collapse [hypersd]. More fundamentally, instead of viewing DMD and RL as two disparate losses added heuristically, we reframe the joint objective as maximizing reward under a distribution-level structural constraint. As the joint gradient can be decomposed into two orthogonal objectives [rlhf, score-base-diff, ppo]: a ”reward ascent” direction that increases human preference, and a ”manifold projection” term that corrects off-manifold displacement by pulling the student back toward the teacher’s regions.The DMD gradient $\nabla_{\!\theta}\mathcal{L}_{\text{dmd}}$ corresponds exactly to the maximum likelihood update projecting the student output distribution toward the learned data distribution of the multi-step teacher $P_{\text{real}}$. When combined linearly with the reward gradient $\nabla_{\!\theta}\mathcal{L}_{\text{rl}}$, any displacement induced by $r$ that exits the teacher’s high-density manifold generates a large discrepancy $s_{\text{real}}-s_{\text{fake}}$. The $\mathcal{L}_{\text{dmd}}$ gradient thus projects the reward ascent strictly along the teacher’s support.
Simultaneously, from the perspective of distillation, pure DMD forces the student to uniformly mimic the teacher, implicitly establishing a ”performance ceiling” capped by distilling $P_{\text{real}}$ equally. Incorporating the RL objective addresses this fundamental limitation: RL steers the distillation toward high-reward regions, allowing the student to surpass the teacher while the DMD regularizer ensures semantic integrity is preserved.
Figure 3: Verification of the joint optimization synergy. Directly conduct RL (red) on the distilled model exhibits high training rewards but suffers from reward hacking and semantic collapse (low DPG score). In contrast, the joint DMD+RL paradigm (green) effectively boosts the performance (better than the DMD baseline (blue) in all metric) while preventing reward hacking. Better to zoom in to check the effect.
To validate the analysis, we conduct a comparative study using SD3-M [sd3] as the base model, with rewards provided by HPS v2.1 [hpsv2] and ReFL-based RL [refl] for training. We compare three paradigms: (i) DMD Only (Blue line); (ii) DMD then RL (Sequential: distillation followed by RL in red line); and (iii) DMD then DMD+RL (Joint approach: distillation followed by joint optimization in green line). We evaluate these models across training/test reward curves and the DPG_Bench score [dpg]. As shown in Fig. 3(a), when RL is introduced (after 1000 iterations), the sequential approach achieves the highest training reward. However, this gain is obtained through hacking score. Fig. 3(b) and (c) reveal that this it suffers from severe reward hacking and catastrophic forgetting. While its training reward climbs, its test reward plateaus, and its DPG score collapses. This confirms that without the strong distributional anchor of the teacher, the few-step model quickly drifts into narrow, high-reward modes that sacrifice semantic integrity [hypersd]. In contrast, the joint optimization demonstrates a clear synergistic effect. First, it effectively breaks the ”performance ceiling” of the teacher; the test reward significantly surpasses the DMD-only baseline. Second, by treating the DMD loss as a persistent structural regularizer, the model keeps improving DPG score while increasing aesthetic quality. This indicates that the DMD objective prevents the RL process from hacking, while the RL objective guides the distillation toward more human-preferred regions.
Table 1: Comparison of different training strategies. The joint approach avoids the ”distillation gap” of ”RL then distillation” pipeline, achieving the highest preference score with $\sim 10\times$ rollout speedup.
Method HPS DPG Time
SD3-M, resolution $\times1024$, local batchsize 16 on H100
Multi-Step RL 31.06 86.43 9.24s $\sim$ 10.57s
Multi-Step RL then Distillation 30.42 85.92 -
Few-Step Joint RL and DMD 31.42 86.08 0.43s $\sim$ 1.21s
Based on the empirical synergy observed above, we formally define the optimization objective. The total loss is formulated as:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{dmd}}+\lambda_{\text{rl}}\mathcal{L}_{\text{rl}},$ (3)
where $\mathcal{L}_{\text{dmd}}$ is the original distribution matching loss with the gradient formulated in Eq. 1, and $\mathcal{L}_{\text{rl}}$ is the plug-and-play reward maximization loss from the RL branch (Eq. 1), the gradient can be differentiable type like ReFL, or policy type like GRPO, etc. The hyperparameter $\lambda_{\text{rl}}$ is a balancing coefficient used to trade off the two losses.
2.4 Reward-Tilted Distribution Matching for Integrating RL into the Early Phase of DMD
While we have demonstrated that DMD and RL can be trained jointly with mutual benefit, we still rely on a two-stage paradigm that incorporates RL only after an initial ”cold start” distillation phase [hypersd, diffinstruct++]. This requirement stems from the observation that reward models typically necessitate a baseline generation capability to produce reliable feedback [refl, srpo]. Specifically, in DMD frameworks with flow matching model, the image used for gradient computation is defined as $\bm{x}_{t}-(0-t)G_{\theta}(\bm{x}_{t})$, where $\bm{x}_{t}$ is iteratively sampled. Notably, t can be any training time-step; thus, when training a 4-step model, the clean image can be generated in only one to four steps. In the early training phase, however, such few-step sampling often fails to produce coherent images, rendering reward signals noisy and unreliable for direct optimization.
Figure 4: Effectiveness of RT-DM and two-stage training strategy. In Stage 1, the RT-DM paradigm (red/green) significantly accelerates preference alignment compared to the DMD baseline (blue). In Stage 2, while ”RT-DM only” (red) eventually saturates due to the constraints of the teacher’s manifold, the transition to joint DMD+RL (green) successfully breaks this performance plateau by directly optimizing the reward score. This complete two-stage approach achieves the highest test rewards and DPG scores. Better to zoom in to check the effect.
To address this, we propose a Reward-Tilted Distribution Matching (RT-DM) paradigm to ”softly” integrate reward signals into the distillation process. Specifically, we utilize the reward score as a weighting factor for the DMD loss. This formulation enables a more flexible optimization objective compared to direct reward maximization. This is because the reward score serves as a static weighting factor, we avoid the necessity of backpropagating through the multi-step denoising process, which allows us to evaluate the reward on the complete trajectory of the few-step model (e.g., 4 steps) without the computational overhead of a large gradient graph, bypassing the need for single-step $\bm{x}_{0}$ approximations to make the reward score more precise and highly valuable as a reference for optimization efforts. More profoundly, we are changing which distribution the student learns by smoothly integrating reward signals to make the student not learn toward the vanilla $P_{\text{real}}$, but toward a reward-tilted teacher distribution defined proportionally to $r(x_{0})p_{\text{real}}$. The gradient is formulated as:
$\nabla_{\!\theta}\,\mathcal{L}_{\text{RT-DM}}$ $=\mathbb{E}_{t}\!\left[\nabla_{\!\theta}\operatorname{KL}(p_{\text{fake},t}\|(p_{\text{real},t})r(\bm{x}_{0}))\right]$ (4)
$=-\mathbb{E}_{t}\!\left[\int\!e^{\frac{r(\bm{x}_{0})}{\left\|r(\bm{x}_{0})\right\|}}\Bigl(s_{\text{real}}\bigl(F_{t})\bigr)-s_{\text{fake}}\bigl(F_{t}\bigr)\Bigr)\frac{\mathrm{d}G_{\theta}(z)}{\mathrm{d}\theta}\,\mathrm{d}z\right]\!.$
To evaluate the effectiveness of the proposed soft integration, we compare three configurations in Fig. 4: (i) DMD Only (Blue); (ii) RT-DM for the entire duration (Red); and (iii) RT-DM followed by joint DMD+RL (Green). As shown in Fig. 4(a) and (b), the RT-DM paradigm (Red/Green) provides a significant ”warm start” effect, accelerating preference alignment in the early phase (Stage 1) compared to the DMD-only baseline. This confirms that weighting the DMD loss with reward scores successfully incorporates preference signals even when few-step generations are still maturing, avoiding the noise instability of direct RL. However, the performance of the ”RT-DM only” approach (Red) eventually saturates. We hypothesize that since RT-DM primarily acts by re-weighting the teacher’s distribution gradients, its ability to drive the model toward high-reward regions is inherently constrained by the teacher’s manifold. To further unlock the model’s potential, we introduce a more direct reward maximization phase (Green) in Stage 2. By transitioning to joint DMD+RL optimization once the model has reached a stable baseline, we can leverage direct RL gradients to aggressively push the performance beyond the initial plateau. As a result, the joint approach achieves the highest test rewards and DPG scores, demonstrating that RT-DM serves as an ideal bridge to transition from pure distillation to direct preference alignment.
2.5 Annealed Distributional Overlap Maximization for Better Initial Distillation
During the initial phase of distillation, DMD encounters a classic ”cold start” challenge: the synthetic samples generated by the few-step student exhibit substantial divergence from the real-world pre-training samples. This profound disjointness between $P_{\text{fake}}$ and $P_{\text{real}}$ impedes the teacher model’s ability to estimate reliable scores, rendering the gradient $\nabla\log P_{\text{real}}-\nabla\log P_{\text{fake}}$ unreliable for optimization. To mitigate this, we introduce a unified principle of annealed distributional overlap maximization. We propose two complementary techniques to artificially increase the overlap between $P_{\text{fake}}$ and $P_{\text{real}}$ early in training, progressively relaxing them as the model converges.
Dynamic Distribution Guidance (DynaDG). DynaDG maximizes overlap primarily by actively shifting the target distribution. The poor initial Image quality causes $P_{\text{fake}}$ to be disjoint from $P_{\text{real}}$ (Figure 6, left). To solve this, we inject trainable LoRA modules [lora] into the real score estimator to ”pull” the perceived $P_{\text{real}}$ toward the nascent $P_{\text{fake}}$ (Figure 6, right). This ensures a reliable optimization signal by bridging the manifold gap. As training progresses, we anneal the LoRA scale in the teacher model toward zero, allowing the real score estimator to seamlessly revert to the true $P_{\text{real}}$ once $P_{\text{fake}}$ is safely anchored.
Figure 5: Impact of dynamic training strategies. (a) Qualitative comparisons show that our dynamic strategies facilitate faster global structure building (highlighted in red boxes). (b) Quantitative DPG scores further confirm that DynaDG and DynaRS provide more reliable optimization signals, consistently outperforming the baseline throughout Stage 1. Better to zoom in to check the effect.
Dynamic Renoise Sampling (DynaRS). While DynaDG moves the distributions, DynaRS maximizes overlap by exploiting regions where distributions naturally intersect. By heavily biasing the sampling of renoise levels $t$ toward higher noise values at the start of training, we ensure that both initial samples and reference samples are dominated by Gaussian noise, placing them in a similar regime. At these high noise levels, the score estimator provides highly reliable gradients that emphasize global structure [lsast, prospect]. As the student generator improves, the renoise bias is progressively annealed to uniform sampling, allowing the model to capture fine-grained textural details from a wider range of noise levels.
Figure 6: Illustration for Dynamic Distribution Guidance.
As shown in Figure 5 (b), our dynamic training strategies significantly enhance the initial training phase. Qualitative results presented in Figure 5 left further corroborate these findings, showing a faster global structure building when DynaDG and DynaRS are employed.
3 Experiments
3.1 Experimental Setup
For training, we adopt denoising-based model (SDXL-Base [sdxl]) and flow-based models (SD3-Medium [sd3], SD3.5-Large [sd35]) for distillation, using prompts from t2i-2M. Additionally, we use ReFL [refl] with DFN-CLIP [dfnclip] and HPSv2.1 [hpsv2] as reward models by default. As for evaluation, we validate the effect of DMDR through extensive experiments. First, we use prompts sampled from a recent public dataset ShareGPT-4o-Image [sharegpt4o-image] to generate images and follow previous work [dmdx, tdm] that report CLIP Score [clipscore], Aesthetic Score [laion5b], Pick Score [pickscore], and Human Preference (HP) Score [hpsv2]. Next, in order to obtain a more comprehensive evaluation, we also compare our distilled models with their teachers on two popular benchmarks, DPG_Bench [dpg] and GenEval [geneval]. Noted that there are some differences from the settings we made in the Method section. We train 1.5K steps in stage1 and stage 2, respectively, to achieve better performance.
Table 2: System-level comparison against state-of-the-art methods. denotes our reproduced results based on the official code. Best performance is marked in Bold.
[…]
SDXL-Base
Base-Model 50 - 34.7588 5.6480 22.1085 27.1477
ReFL [refl] 50 RL 35.2107 5.9045 22.8076 33.1874
LCM [lcm] 2 Distill 28.4664 5.1026 20.0603 17.6837
Lightning [sdxl-lightning] 1 Distill 32.0283 5.6761 21.4868 26.3615
DMD2 [dmd2] 1 Distill 34.3046 5.6238 21.7293 26.5986
DMDR (ours) 1 Distill & RL 35.6241 6.1184 22.8498 32.0065
DMD2 [dmd2] 4 Distill 34.5169 5.7043 22.1546 28.5655
DMD2-PSO [pso] 4 Distill then RL 34.0128 5.8032 22.2644 29.8732
DMDR (ours) 4 Distill & RL 35.4940 6.0324 22.7122 32.9832
SD3-Medium
Base-Model 50 - 34.9025 5.5942 22.1801 28.4021
ReFL [refl] 50 RL 35.1851 5.7821 22.0064 32.0745
DMD2 [dmd2] 4 Distill 33.9421 5.6137 21.7688 27.3675
Flash [flashdiff] 4 Distill 34.2634 5.6702 21.5921 26.6542
TDM [tdm] 4 Distill 34.0301 5.6250 22.0010 27.7522
Hyper-SD [hypersd] 8 Distill then RL 32.0234 5.2489 20.2831 22.4544
DMDR (ours) 4 Distill & RL 35.0142 5.8876 22.4892 32.1145
SD3.5-Large
Base-Model 50 - 35.5509 5.7014 22.4856 28.8135
LADD [ladd] 4 Distill then RL 35.0480 5.4514 22.2451 27.8470
DMDR (ours) 4 Distill & RL 35.9757 6.1541 22.9072 32.7368
Method for comparison. We compare our method against several categories of existing work, including foundational multi-step base models [sdxl, sd3, sd35], RL only approaches [refl], distillation only methods [lcm, sdxl-lightning, dmd2, flashdiff, tdm], distillation then RL frameworks [hypersd, pso]. We do not compare with methods like Flash-DMD [flash-dmd] and Diff-Instruct++ [diffinstruct++] that do not provide model weights or open source training code because we cannot guarantee a fair and reproducible evaluation.
3.2 Main Results
We begin by assessing DMDR’s text-to-image generation performance using prompts sampled from ShareGPT-4o-Image [sharegpt4o-image], ensuring that there is no overlap with the prompts utilized during training. Table 3.1 presents a summary of the distillation results in comparison to other methods. Models distilled through our approach exhibit state-of-the-art (SOTA) capabilities in terms of prompt coherence and the generation of high-quality, aesthetically pleasing images. Notably, our method demonstrates effectiveness across various architectures (e.g., UNet, Transformer), model sizes (e.g., 2B, 8B), and paradigms (e.g., flow-based, denoising-based), highlighting the broad applicability and universality of the proposed approach. Additionally, we provide qualitative comparisons in Figure 7, where our method produces images characterized by superior quality and enhanced prompt coherence.
Table 3: GenEval Comparison: 4-step (ours) $vs.$ its multi-step teacher.
Model Overall Single Two Count. Colors Pos. Attr.
SDXL-Base
Teacher 0.55 0.98 0.74 0.39 0.85 0.15 0.23
Ours 0.57 0.99 0.76 0.44 0.85 0.12 0.25
SD3-Medium
Teacher 0.62 0.98 0.74 0.63 0.67 0.34 0.36
Ours 0.65 0.99 0.84 0.57 0.81 0.27 0.45
SD3.5-Large
Teacher 0.71 0.98 0.89 0.73 0.83 0.34 0.47
Ours 0.72 0.99 0.93 0.69 0.80 0.31 0.59
Table 4: DPG_Bench Comparison: 4-step (ours) $vs.$ its multi-step teacher.
Model Overall Global Entity Attribute Relation Other
SDXL-Base
Teacher 74.65 83.27 82.43 80.91 86.76 80.41
Ours 76.48 83.72 82.58 83.69 84.76 83.54
SD3-Medium
Teacher 84.08 87.90 91.01 88.83 80.70 88.68
Ours 85.30 90.49 90.07 90.56 87.17 91.23
SD3.5-Large
Teacher 84.12 91.48 90.22 87.81 91.20 89.49
Ours 85.33 90.47 90.53 90.68 87.44 90.20
Figure 7: Visual comparison between the teachers, selected competing methods [dmd2, hypersd, ladd], and ours. All images are generated using identical noise. Our model produces images with superior quality and prompt coherence.
To further verify our method’s effectiveness and provide a more comprehensive comparison between the few-step distillation model and the multi-step teacher model, we conduct the comparison on two commonly used benchmarks [dpg, geneval], Although we do not perform RL optimization on the metric of these two benchmarks, table 4 and  4 show that the overall score of our four-step model consistently outperforms its multi-step teacher on DPG_Bench and GenEvalacross all base models. These results indicate that DMDR has successfully freed the student model from the constraints of the multi-step teacher model and stimulated its capabilities during the distillation process.
Table 5: Ablation on different components. Default settings are marked in purple.
(a) Ablation on different reward models.
Method CLIP Aesthetic Pick HP DPG
Score Score Score Score overall
Pick [pickscore] 34.3403 5.9021 22.9674 31.7665 84.86
AE [laion5b] 34.1748 6.1024 22.3858 31.6620 84.61
CLIP [clipscore] 35.2087 5.5832 22.0644 28.0507 85.10
HPS [hpsv2] 34.4088 5.9021 22.5108 32.9065 85.04
CLIP + HPS 35.0142 5.8876 22.4892 32.1145 85.30
(b) Ablation on RL algorithm in stage 2.
Method CLIP Aesthetic Pick HP DPG
Score Score Score Score overall
stage 1 init 33.3648 5.5744 21.0070 28.8371 82.4371
w/ RL (ReFL) 35.0142 5.8876 22.4892 32.1145 85.30
w/ RL (DPO) 34.0234 5.8340 21.9844 30.8352 85.00
w/ RL (GRPO) 34.6675 5.9324 22.3248 31.4474 85.08
(c) Ablation on Value of $\lambda_{\text{rl}}$.
Value CLIP Aesthetic Pick HP DPG
Score Score Score Score overall
0.1 34.7422 5.8508 22.3064 31.6565 85.24
0.5 35.0142 5.8876 22.4892 32.1145 85.30
1.0 35.0087 5.84430 22.4709 32.8982 85.04
2.0 35.0322 5.7806 22.1098 32.7683 82.48
3.3 Ablation Study
In this section, we conduct a systematic investigation into the key components of DMDR. Unless otherwise specified, all experiments use the SD3-M base model with a 4-step sampling configuration.
Versatility across reward models. A key advantage of ourusing a DINOv2-based reward model to maximize classification accuracyDMDR is the ability to steer the distillation process toward diverse human preferences. As shown in Table 5(a), we evaluate the framework using different reward models (RMs) targeting specific attributes. The results provide a clear insight: optimizing with a specific RM leads to a significant performance gain in its corresponding metric, confirming that DMDR effectively injects the desired preference signal into the few-step manifold. Furthermore, by integrating multiple rewards (CLIP + HPS), the model achieves a more balanced and superior generative performance, similar to the findings in DanceGRPO [dancegrpo]. This flexibility allows users to customize the distillation ”flavor” according to specific downstream tasks.
Compatibility with RL paradigms. To verify the algorithmic universality of our unified framework, we compare three distinct RL strategies in the Stage 2 joint optimization phase: ReFL [refl], DPO [diffusiondpo, pso], and GRPO [dancegrpo, flowgrpo]. As shown in Table 5(b), all three algorithms yield substantial improvements over the ”Stage 1 init” baseline across all evaluation dimensions. Furthermore, an interesting observation is that using the ReFL in DMDR achieves overall better performance, and our analysis is that: traditional ReFL for multi-step models ignores earlier sampling steps and only trains the last few steps before the output image [refl] because the large computational graph generated by the multi-step denoising process would result in a huge memory consumption. Subsequent solutions [srpo, deepreward] mainly focus on saving forward noise trajectory and sampling a subset of steps to optimize, although this can allow the gradient to return to the early denoising step, it is also prone to error accumulation. On the contrary, in our few-step model, the model is trained to predict a clean image at each step’s state [dmd2], which enables the gradient of ReFL to directly return to the initial state, thus obtaining an effective optimization.
The balancing Role of DMD Regularization. The synergy between distillation and RL is governed by the balancing coefficient $\lambda_{\text{rl}}$ in our joint objective. Table 5(c) illustrates the trade-off between reward maximization and distributional stability. When $\lambda_{\text{rl}}$ is set too low; the model remains constrained by the teacher’s manifold, limiting the potential for preference alignment. Conversely, as it increases beyond an optimal threshold, we observe a noticeable decline in a general capacity (reflected by the DPG overall score), despite high reward scores. This trend provides a crucial insight: the DMD loss serves as a vital structural regularizer. It prevents the model from ”reward hacking”(We also showed in Figure 3. The optimal balance ensures that the model pushes the boundaries of quality while staying anchored to the teacher’s robust generative priors.
4 Limitation and Discussion
Quality vs. Diversity. To further analyze the behavior of DMDR, we conduct class-conditional experiments on ImageNet $\times256$ using the SiT [sit] backbone with representation alignment pretraining ckpt [repa, sra] as initialization for 1-step generation. As shown in Table 6, a fundamental trade-off between generative quality and diversity emerges. Our results reveal that both Classifier-Free Guidance (CFG) (this also observed in Decoupled-DMD [decoupled-dmd]) and Reinforcement Learning (here we use a DINOv2-based [dinov2] reward model to maximize classification accuracy) act as ”density sharpeners” during the distillation process. This phenomenon suggests that DMDR effectively ”sacrifices” the teacher’s distributional uncertainty in exchange for the stability and quality required for few-step sampling. We believe this reduction in diversity is not a limitation unique to our approach, but a characteristic of score-based distillation and preference alignment, which steer the model toward high-reward, high-confidence regions. This trade-off is often desirable in practical few-step applications where sample excellence and prompt adherence are prioritized over exhaustive distribution coverage. But at the same time, further investigation into finding the optimal balance between quality and diversity is left for future work.
Table 6: 1-step SiT DMDR Results on ImageNet $256\times 256$. We report the performance with different CFG scales and RL optimization. Metrics include FID [fid], Recall [pre-rec], Inception Score (IS) [is], and Precision [pre-rec].
CFG RL FID $\downarrow$ Recall $\uparrow$ Inception Score $\uparrow$ Precision $\uparrow$
No No 2.13 0.64 232.15 0.77
1.5 No 5.20 0.48 387.24 0.88
4 No 15.70 0.11 453.65 0.86
No Yes 6.95 0.40 416.01 0.90
1.5 Yes 13.43 0.24 494.43 0.93
5 Related Work
Distribution matching distillation. Distribution Matching Distillation [dmd] (DMD) is the first work to successfully apply the principle score-based distillation [score-base-diff, prolificdreamer, dreamfusion] to large-scale text-to-image models. Intuitively, it strives to ensure that any sample realized by the student at a given noise level occurs with exactly the same probability as it would under the teacher’s distribution, thereby preserving the multi-step generative priors in the few-step model. From then on, a lot of follow-up work emerged [dmd2, dmdx, tdm, sd3.5flash, f-distill, flash-dmd, twinflow], for example, DMD2 [dmd2] employs a discriminator to align the student model distribution with a specific target distribution like GAN [gan] does; TDM [tdm] incorporates DMD loss in the sampling process of the student model for better alignment; ff–distill [f-distill] replaces the original reverse Kullback–Leibler (KL) divergence to the proposed $f$-divergence for covering different divergences with different properties. Our work also starts with DMD, but we don’t focus on how to better “imitate” the teacher. Instead, we aim to incorporate reinforcement learning in the distillation process to enable a more controllable and preference-aligned distillation.
Reinforce learning for diffusion models. Reinforce learning helps to align diffusion models to human preferences by training a reward model and using it to guide generation [ddpo, flowgrpo, refl, srpo, diffusiondpo, dancegrpo, diffusionnft]. There are many algorithms to conduct RL, for example, DDPO [ddpo] adapts PPO via image log-likelihoods; ReFL [refl] bypasses likelihoods by optimizing outputs with frozen-reward gradients; Diffusion-DPO [diffusiondpo] adapts DPO to diffusion for paired human preference data; and recent GRPO extensions [flowgrpo, dancegrpo] use GRPO to diffusion models by coupling the training loss with SDE samplers. Although significant progress has been made in RL for diffusion models, most of the work has focused on multi-step models. Here, we find that these algorithms are also adapted to the few-step model when carried out in conjunction with distribution matching distillation.
Trajectory-based distillation. Trajectory-based distillation [lcm, hypersd, rcm, piflow], which typically aims to simulate teacher ODE trajectories on the instance level. Early work [kdgen] regresses the teacher’s ODE integral in one step, producing blurry $\ell_{2}$$x_{0}$ estimates, and thus suffers from degraded quality. Progressive distillation [hypersd, shortcut-model] mitigates this by a multi-stage pipeline that enlarges the student’s step size and halves its NFE each stage by distilling the prior stage’s trajectory into fewer steps, However, this not only makes the training a multi-stage process (e.g., NFE from 50 to 25 then to 10, etc.), but also leads to cumulative errors. Meanwhile, consistency models [lcm, scm] aim to learn a consistency function that maps the point at an arbitrary time t on the teacher’s PF-ODE trajectory to the initial point. However, the student model must be constructed implicitly using either inaccurate finite differences or expensive Jacobian–vector products (JVPs) and the quality is still limited due to the accumulation of errors into the integrated state, which limits the application in large-scale scenarios [rcm, piflow].
Adversarial distillation. Adversarial distillation [add, apt, ladd, apt2] can be regarded as another form of distribution matching distillation. The difference is that adversarial distillation aim to estimate the distribution of both the student model and the real data or the teacher model through a discriminator model [gan] to match the distribution. Representative works like Adversarial Diffusion Distillation (ADD) [add] force the few-step student diffusion model to fool a discriminator which is trained to distinguish the generated samples from real images. Follow-up works [apt, apt2] applied this idea to other fields. However, such a training method can cause many instabilities, similar to traditional GANs [gan, improvedgan], hence requires a lot of tricks to stabilize training [animatediff-l, r1reg, train-gan].
Other distillation works. Other distillation works mainly focus on combining the distillation at the distribution level and the trajectory level [sana-sprint, rcm, tdm, pose]. For example, SANA-Sprint [sana-sprint] combines the ideas of LADD [ladd] and sCM [scm] and achieves fast convergence and high fidelity generation while retaining the alignment advantages of sCMs. TDM [tdm] and rCM [rcm] combine DMD [dmd] and LCM [lcm] together to remedy the quality issues of LCM.
Reward models for diffusion post-train. Unlike Large Language Models (LLMs), where reinforcement learning (RL) can often leverage verifiable, rule-based rewards [deepseekmath, deepseekr1, glm] (e.g., code execution success or mathematical correctness), diffusion-based image generation is inherently subjective. The evaluation of generated images relies heavily on multifaceted criteria such as aesthetic quality, instruction following, and so on, which are difficult to quantify through rigid, rule-based functions. Consequently, the development of robust reward models (RMs) has become a critical prerequisite for effective RL-based alignment in diffusion models [refl, rewarddance, clipscore, hps, visionreward, q-align]. Early methods, such as CLIP-score [clipscore], primarily focus on measuring the semantic alignment between the text prompt and the generated image. To address the need for visual fidelity and artistic appeal, metrics like Aesthetic Score [laion5b] and HPS (Human Preference Score) [hps, hpsv2, hpsv3] have been introduced to capture the nuance of human aesthetic judgment. More recently, the reward models become unified [unified-reward, visionreward, unified-p-reward] and aggregate multi-dimensional preferences—ranging from low-level texture details to high-level semantic instruction following—into a unified scoring function. While our primary contribution lies in proposing a novel algorithm that integrates few-step distillation with reinforcement learning, our empirical results highlight the pivotal role of the reward model in our pipeline.
6 Conclusion
We introduces DMDR, a unified framework that transforms Distribution Matching Distillation (DMD) and Reinforcement Learning (RL) from independent, sequential stages into a synergistic training pipeline. We demonstrate that joint optimization is mutually beneficial: RL steers distillation toward high-reward regions to break the ”teacher ceiling,” while DMD regularizes RL to effectively mitigate reward hacking. By employing a two-stage strategy—incorporating RT-DM and dynamic training followed by joint optimization—DMDR achieves state-of-the-art few-step performance across diverse architectures. Ultimately, our approach enables few-step models to not only match but consistently surpass the visual quality and prompt adherence of their original multi-step teacher models.